# Relevance-driven Decision Making for Safer and More Efficient Human Robot Collaboration

### v1

@@ Abstract
Human intelligence possesses the ability to effectively focus on important environmental components, which enhances perception, learning, reasoning, and decision-making. Inspired by this cognitive mechanism, we introduced a novel concept termed relevance for Human-Robot Collaboration (HRC). Relevance is defined as the importance of the objects based on the applicability and pertinence of the objects for the human objective or other factors. In this paper, we further developed a novel two-loop framework integrating real-time and asynchronous processing to quantify relevance and apply relevance for safer and more efficient HRC. The asynchronous loop leverages the world knowledge from an LLM and quantifies relevance, and the real-time loop executes scene understanding, human intent prediction, and decision-making based on relevance. In decision making, we proposed and developed a human robot task allocation method based on relevance and a novel motion generation and collision avoidance methodology considering the prediction of human trajectory. Simulations and experiments show that our methodology for relevance quantification can accurately and robustly predict the human objective and relevance, with an average accuracy of up to 0.90 for objective prediction and up to 0.96 for relevance prediction. Moreover, our motion generation methodology reduces collision cases by 63.76% and collision frames by 44.74% when compared with a state-of-the-art (SOTA) collision avoidance method. Our framework and methodologies, with relevance, guide the robot on how to best assist humans and generate safer and more efficient actions for HRC.
@@ I Introduction
Robots and automation systems are becoming increasingly critical in enhancing productivity, efficiency, and precision in industry [1, 2, 3] and daily human life [4, 5]. However, they still struggle to match the exceptional cognitive capabilities of human beings. Among all cognitive mechanisms in human brains, humans are extraordinary at selectively focusing on important components within the environment, which guide our perception, learning, reasoning, and decision-making in daily life. It enables us to focus on relevant stimuli while filtering out irrelevant information in our environment, thus playing a crucial role in more efficient spatial perception, scene understanding, and decision-making [6, 7].
To empower the robot with similar capabilities, in [8], we defined a new and important concept and scene understanding approach, termed relevance, and developed methodologies for quantifying relevance with a novel event-based framework and a probabilistic methodology based on a new scene representation. Relevance is defined as the pertinence of objects in the scene based on the human objective or other factors. An example of relevance is shown in Fig. 1.
Figure 1: Illustration of an example of relevance in a tabletop setup. The human is currently cutting an orange. The robot needs to reason about the scene, the human, and other factors to determine the human’s objective. With the objective and possible tasks, relevant classes and elements of objects are determined for how to best assist the human. With the quantified relevance, the robot is informed about which objects to fetch and deliver for the best human assistance and which objects to focus on in other functions to improve performance.
@@
Relevance offers three key benefits in human-robot collaboration (HRC). First, accurately determined relevance enables the robot to better understand object applicability to the scene and potential interaction sequences, enhancing the efficiency, safety, and fluency of HRC. Second, focusing on relevant objects allows the robot to optimize its computational resources, improving speed and safety [9]. Third, relevance unites human, task, and scene models, facilitating more accurate predictions and reasoning by leveraging improvements from various areas.
@@
Building on the work in [8], this paper further develops and presents a framework for determining relevance and applying it to a novel real-time decision-making component that enhances safety in HRC. In this framework, two loops are running asynchronously, namely the async loop and the real-time loop. The real-time loop performs continuous real-time computation, including scene understanding, human intent prediction, and decision-making. The asynchronous loop leverages the world knowledge from Large Language Models (LLM) to predict the potential human objective and relevance. The real-time loop utilizes those results for decision-making to achieve proactive and safe HRC. Although the loop operates asynchronously, its execution rate dynamically adjusts based on the scene’s conditions.
@@
In the decision-making module, we developed a real-time methodology with human-robot task allocation and a novel motion planning methodology. With relevance, the traditional Artificial Potential Field (APF) methods for motion planning and collision avoidance are enhanced by a virtual obstacle constructed based on the future projection of the human motion and associated repulsive force formulation. In this way, the robot can behave in a proactive manner for safer and more efficient HRC. Simulation results demonstrate that our decision making module reduces the collision cases and frames by 63.76% and 44.74 %, respectively.
@@
In summary, the paper’s contributions are as follows: (1) We proposed and developed a novel methodology to quantify relevance with human objective prediction and relevant object prediction based on LLMs. (2) We proposed and developed a novel two-loop framework to integrate LLM inference and relevance quantification into a real-time application by leveraging the asynchronous frameworks based on different inference requirements of different components. (3) We proposed and developed a novel methodology of decision making based on relevance. Simulations and experiments validate enhanced safety and lower collision probability with our Relevance-based Artificial Potential Field (RAPF).
@@ II-A Saliency and Attention
Relevance fundamentally differs from previous works that aim to identify and underline the important regions in the scene, such as attention and saliency. Saliency focuses on identifying the visually prominent and conspicuous features in the image [10, 11, 12], while attention focuses on current short-term, reactive, and simple tasks without considering the future projection [13, 14]. Relevance fundamentally differs from those works in that relevance represents the inherent relationship between objects because of the human objective and the applicability of objects to achive the objective. Relevance considers the interconnection among environmental understanding, human models, task models, etc.
II-B LLMs for Robotics
Large Language Models (LLMs) are revolutionizing robotics by enabling multi-modality reasoning [15], general prediction without transfer learning [16], and flexible connections between modules in robotic frameworks [17]. In this paper, we utilized LLMs for a dramatically different and novel application, i.e., relevance determination. Moreover, the current inference time of LLMs is longer than that of real-time computation requirements. Existing works with LLMs in robotics mainly apply to non-real-time policy generation in a static or pre-defined environment. Our unique two-loop design enables knowledge retrieval of LLMs in real-time applications by leveraging asynchronous computation.
@@ III-A Problem definition
Let $S=\{C_{1},C_{2},...,C_{n}\}$ be a set of class of objects in a visual scene, where each $C_{i}$ represents a class of objects and $n$ is the number of classes of objects in the scene. The set of relevant elements $\mathcal{E}_{r}$, which contains all the relevant elements to the human’s objective and associated tasks, is predicted and determined with the available information. Based on $\mathcal{E}_{r}$, the robot generates actions to assist the human to achieve the human’s objective.
@@ III-B Framework overview
An overview of the framework is shown in Fig. 2. Our framework for relevance quantification and safer HRC consists of two loops. The first loop fulfills the function of perceiving the world, understanding the scene, and decision making based on relevance for safer HRC. The second loop is an asynchronous loop that leverages LLMs and quantifies relevance. By optimally coordinating the two loops, the world knowledge of the LLM can be incorporated into real-time operation of the robots.
Figure 2: The overview of the methodology for relevance quantification and application of relevance for safer human-robot collaboration. The inputs to the algorithm pipeline are the stream images from the camera, which are processed with scene understanding algorithms. Based on the scene understanding, the human intent is classified and fed into the LLM in an asynchronous manner. With the LLM results, the relevance can be quantified, which informs the decision-making module to generate actions for the robot for safer HRC.
@@ III-C Scene understanding
The input to the algorithm is the stream of observations from the robot sensors. The images are processed with scene understanding algorithms, such as YOLO [20] for object detection, MMPose [21] for human pose estimation, CLIP [22] for scene grounding, etc.
@@ III-D Human intention prediction
The module of human intention prediction determines current human actions and aggregates them into a history record. Here, we implement an algorithm that leverages Bayesian networks to fuse three modalities of human information: head orientation, hand orientation, and hand movement. By fusing three modalities for human intention recognition, the accuracy and precision of the prediction are dramatically increased because different modality provides valuable information in different stages of the actions.
@@ III-E Relevance Quantification with LLMs
The first step in quantifying relevance with LLM is contextualizing the scene. Currently, this is achieved with a fixed prompt $\mathcal{P}$ designed and optimized specifically for the setup and the problem of HRC. However, in the future, automatic scene contextualization and prompt construction can be considered and developed. In the prompt $\mathcal{P}$, the environment, the objects in the scene, and the action history of the human are automatically generated from the scene understanding and human intention prediction modules.
The outputs of the LLM module are the prediction of the human objective and the set of relevant elements $\mathcal{E}_{r}$. The relevance of each element is quantified to be
$R(e_{i})=\begin{cases}1&\text{if }e_{i}\in\mathcal{E}_{r}\\ 0&\text{otherwise}\end{cases}$ (1)
With this method, world knowledge of an LLM can be extracted for a general environment and setup without any training and transfer learning.
@@ III-F Decision making - task allocation
In the problem definition, we define the subscript $j$ representing the index of elements such that $e_{j}\in\mathcal{E}_{r}$. We define the position vector of element $e_{j}$ in the world coordiante as $\mathbf{p}_{e_{j}}$, the position of the common destination of relevant elements as $\mathbf{p}_{d}$, the initial location of the robot as $\mathbf{p}_{r_{0}}$, the velocity of the robot as $v_{r}$, the velocity of human as $v_{h}$, the start-up delay for the human to finish the current task as $d_{h}$. We consider $d_{h}$ because there is a temporal delay between the time the robot and the human start to fulfill the allocated tasks. We further define $T_{r}$ and $T_{h}$ as the time durations for the robot or the human to finish the assigned tasks, which can be computed as:
@@ III-G Decision making - motion planning
where $A$ is the magnitude of the attractive force, $d_{g}$ is the distance between the robot and its goal, $\alpha_{a}$ is the constant that controls how the attractive force decreases with distance, and $\mathbf{\hat{u}_{a}}$ is the unit vector from the robot to the goal.
@@
To consider the future motion of the human hand, we developed a novel methodology to construct an ellipsoid virtual obstacle as shown in Fig. 3. The two endpoints of the major axis of the ellipsoid are the current position of the human hand $\mathbf{p}_{h}$ and the position of the predicted hand destination $\mathbf{p}_{t}$. The semi-axes $a$, $b$, and $c$ of the ellipsoid are given by:
@@
where $\mathbf{\hat{u}_{h\rightarrow t}}$ represents the unit vector from $\mathbf{p}_{h}$ to $\mathbf{p}_{t}$, and $t_{s}$ is a time factor we proposed in the unit of $s$ reflecting the available safety buffer in terms of time. $k_{r}$ decreases as $\mathbf{p}_{c}$ becomes farther away from $\mathbf{p}_{h}$. Without $k_{r}$, $\mathbf{F}_{r_{v}}$ will consider the virtual obstacle too early and conservatively, which will result in unnecessary detour and a longer robot trajectory.
@@
This total force determines the direction and magnitude of the robot’s motion, allowing it to avoid both physical and virtual obstacles while moving towards its goal.
@@ IV-A Dataset
The performance of human objective prediction and relevance prediction is assessed via the Breakfast Actions Dataset [24], which comprises a variety of typical human activities performed during breakfast time (e.g., preparing coffee, cooking pancakes, and making hot chocolate, etc.). For each test, ground truth data provides the ground truth objective (GTO), and the ground truth plan (GTP), which describes the sequence of actions performed by an individual in the execution of the objective. In our evaluation, a step ratio of segments from the GTP is contextualized and input into LLM to predict the human objective and the set of relevant elements. Three distinct ratio step values: 0.25, 0.5, and 0.75 are employed. The objective prediction is evaluated through a manual assessment process. The predicted relevant objects are evaluated automatically with word matching. The metrics for the evaluation are percentages of tests with correct prediction of the objective and the relevant elements.
@@ IV-B Simulation Development and Setup
To verify the effectiveness of relevance for safer HRC, we developed a simulator as shown in Fig. 4. In the simulator, we constructed the environment with a table (the cyan cuboid) and objects on top of the table. The height of the table is set to be 73 cm. The table size is 180 cm in the $x$ direction, 6 cm in the $z$ direction, and 76 cm in the $y$ direction. The objects on the table consist of two parts: the necessary objects for the objective and randomly added objects from a list of kitchen objects. The locations for each objects on the tabletop are randomly selected from a collection of locations uniformly distributed on a half circle with a radius of 60 cm. The diameter of each object is 8 cm.
@@ V-A Accurate Objective and Relevance Prediction
TABLE I: Performance of objective prediction and relevance prediction for the 10 objectives in the Breakfast Actions Dataset with various step ratios of action as inputs. These results show that our methodology can accurately and robustly predict the human objective and relevance.
@@
We first analyze the performance of objective prediction using our methodology, upon which relevance determination depends. Upon further examination of individual objectives, certain objectives, such as the preparation of cereal, coffee, hot chocolate, juice, salad, and tea, exhibit consistently high predictive accuracy across all evaluated ratio steps. The prediction accuracy for those objectives is high even at a low step ratio of 0.25, demonstrating our methodology can accurately identify the human’s objective at an early stage of the human action and motion. Consistent with expectations, an increase in the “ratio steps” parameter correlates positively with enhanced prediction accuracy. The contextualization of the scene with a higher step ratio will contain more indicative cues about the objectives and thus improve the objective prediction accuracy. The inaccuracy for fried eggs is attributed to the inherent similarity between the objectives involving eggs, such as omelettes and scrambled eggs. However, the relevance for those objectives are very similar.
Next, we analyze the performance of relevance prediction using our methodology based on the predicted objective. The relevance prediction performs extraordinarily for specific objectives, including hot chocolate, juice, and salad, which achieves a 100% accuracy at the step ratio of 0.25. The prediction accuracy of all other objectives achieves at least 0.85 at the step ratio of 0.75. For the relevance prediction, the accuracy also increases with the step ratio. It’s worth noting that the relevance prediction for the objectives predicted with low accuracy can still be accurate. This is attributed to the fact that similar objectives, such as omelettes and scrambled eggs, share the same relevance, which bolster the relevance prediction despite inaccuracies in objective identification.

### v2

@@ Abstract
Human brain possesses the ability to effectively focus on important environmental components, which enhances perception, learning, reasoning, and decision-making. Inspired by this cognitive mechanism, we introduced a novel concept termed relevance for Human-Robot Collaboration (HRC). Relevance is a dimensionality reduction process that incorporates a continuously operating perception module, evaluates cue sufficiency within the scene, and applies a flexible formulation and computation framework. In this paper, we present an enhanced two-loop framework that integrates real-time and asynchronous processing to quantify relevance and leverage it for safer and more efficient human-robot collaboration (HRC). The two-loop framework integrates an asynchronous loop, which leverages an LLM’s world knowledge to quantify relevance, and a real-time loop, which performs scene understanding, human intent prediction, and decision-making based on relevance. HRC decision-making is enhanced by a relevance-based task allocation method, as well as a motion generation and collision avoidance approach that incorporates human trajectory prediction. Simulations and experiments show that our methodology for relevance quantification can accurately and robustly predict the human objective and relevance, with an average accuracy of up to 0.90 for objective prediction and up to 0.96 for relevance prediction. Moreover, our motion generation methodology reduces collision cases by 63.76% and collision frames by 44.74% when compared with a state-of-the-art (SOTA) collision avoidance method. Our framework and methodologies, with relevance, guide the robot on how to best assist humans and generate safer and more efficient actions for HRC.
@@ I Introduction
Robotic and automation systems are becoming increasingly critical in enhancing productivity, efficiency, and precision in industry [1, 2, 3] and daily human life [4, 5]. However, they still struggle to match the exceptional cognitive capabilities of human beings. Among all cognitive mechanisms in human brains, humans are extraordinary at selectively focusing on relevant stimuli while filtering out irrelevant information in our environment, which plays a crucial role in more efficient spatial perception, scene understanding, and decision-making [6, 7]. This mechanism is regulated by a structure of neurons in the brainstem, known as the Reticular Activating System (RAS), which filters sensory information and determines its relevance. The earliest paper on RAS in the human brain can be traced to the work in [8]. It laid the foundation for understanding the RAS and its role in consciousness, attention, and alertness.
To empower robots with similar capabilities, in [9], we defined a new and important concept and dimensionality reduction approach, termed relevance, and developed methodologies for quantifying relevance with a novel event-based framework and a probabilistic methodology based on a new scene representation. Relevance is defined as a dimensionality reduction process that continuously perceives the scene, identifies and detects its elements, and organizes them into a hierarchical representation, such as attribute classes. This process utilizes contextual cues to selectively reduce input dimensionality. When cues are insufficient, relevance-driven processing iteratively gathers additional cues and reapplies dimensionality reduction until sufficient information is obtained.
@@
Relevance offers three key benefits in human-robot collaboration (HRC). First, accurately determined relevance enables the robot to better understand object applicability to the scene and potential interaction sequences, enhancing the efficiency, safety, and fluency of HRC. Second, focusing on relevant objects allows the robot to optimize its computational resources, improving speed and safety [10]. Third, relevance unites human factors, task models, scene relationships, and even information sufficiency, facilitating more accurate predictions and reasoning by leveraging advancements from various areas.
@@
Building on the work in [9], this paper further develops and presents a framework to leverage an AI toolkit to determine relevance and apply it to a novel real-time decision-making component that enhances safety in HRC. In this framework, two loops are running asynchronously, namely the async loop and the real-time loop. The real-time loop performs continuous real-time computation, including scene understanding, human intent prediction, and decision-making. The asynchronous loop leverages the world knowledge from Large Language Models (LLM) in the AI toolkit to predict the potential human objective and relevance. The real-time loop utilizes those results for decision-making to achieve proactive and safe HRC. The execution rates of the two asynchronous loops adjust dynamically based on the scene’s conditions.
@@
In the decision-making module, we developed a real-time methodology with human-robot task allocation and a novel motion planning methodology based on relevance. The traditional Artificial Potential Field (APF) methods for motion planning and collision avoidance are enhanced by a virtual obstacle constructed based on the future projection of the human motion and associated repulsive force formulation. In this way, the robot can behave in a proactive manner for safer and more efficient HRC. Simulation results demonstrate that our decision making module reduces the collision cases and frames by 63.76% and 44.74 %, respectively.
@@
In summary, the paper’s contributions are as follows: (1) We proposed and developed a novel methodology to quantify relevance with human objective prediction and relevant object prediction based on an AI toolkit. (2) We proposed and developed a novel two-loop framework to integrate the LLM inference in the AI toolkit and relevance quantification into a real-time application by leveraging the asynchronous frameworks based on different inference requirements of different components. (3) We proposed and developed a novel methodology of decision-making based on relevance. Simulations and experiments validate enhanced safety and lower collision probability with our Relevance-based Artificial Potential Field (RAPF).
@@ II-A Saliency and Attention
Relevance is a broad concept that covers previous works identifying and highlighting the important features in the scene, such as attention and saliency. Saliency focuses on identifying the visually prominent and conspicuous features in the image [11, 12, 13], while attention focuses on current short-term, reactive, and simple tasks without considering the future projection [14, 15]. Relevance possesses more advanced functionalities by considering information sufficiency, hierarchical scene representation, dynamically scheduled perception modules, etc. Furthermore, relevance is characterized by a more flexible formulation and computational principle than existing methods, which results in unique and additional advantages leading robots toward Artificial General Intelligence (AGI). A more detailed comparison can be found in [9].
II-B LLMs in the AI Toolkit for Robotics
Large Language Models (LLMs) are revolutionizing robotics by enabling multi-modality reasoning [16], general prediction without transfer learning [17], and flexible connections between modules in robotic frameworks [18]. In this paper, we utilized LLMs in the AI toolkit for a dramatically different and novel application, i.e., relevance determination. Moreover, the current inference time of LLMs is longer than that of real-time computation requirements. Existing works with LLMs in robotics mainly apply to non-real-time policy generation in a static or pre-defined environment. Our unique two-loop design enables knowledge retrieval of LLMs in real-time applications by leveraging asynchronous computation.
@@ III-A Problem Definition
Let $S=\{C_{1},C_{2},...,C_{n}\}$ be a set of class of objects in a visual scene, where each $C_{i}$ represents a class of objects and $n$ is the number of classes of objects in the scene. The set of relevant classes $\mathcal{C}_{r}$ and the set of relevant elements $\mathcal{E}_{r}$, based on the human’s objective and associated tasks, is predicted and determined with the available information. Based on $\mathcal{C}_{r}$ and $\mathcal{E}_{r}$, the robot generates actions to assist the human in achieving the human’s objective.
@@ III-B Framework Overview
An overview of the framework is shown in Fig. 1. Our framework for relevance quantification and safer HRC consists of two loops. The first loop fulfills the function of perceiving the environment, understanding the scene, and decision making based on relevance for safer HRC. The second loop is an asynchronous loop that leverages modules in an AI toolkit and quantifies relevance. In this paper, we leverage an LLM for the purpose of illustration. By optimally coordinating the two loops, the world knowledge of the LLM in the AI toolkit can be incorporated into real-time operation of the robots.
Figure 1: The overview of the methodology for relevance quantification and application of relevance for safer human-robot collaboration. The inputs to the algorithm pipeline are the stream images from the camera, which are processed with scene understanding algorithms in the AI toolkit. Based on the scene understanding, the human intent is classified and fed into the LLM within the AI toolkit in an asynchronous manner. With the LLM results, the relevance can be quantified, which informs the decision-making module to generate actions for the robot for safer HRC.
@@ III-C Scene Understanding
The input to the algorithm is the stream of observations from the robot sensors. The images can be processed with scene understanding algorithms in the AI toolkit, such as YOLO [21] for object detection, MMPose [22] for human pose estimation, CLIP [23] for scene grounding, etc.
@@ III-D Human Intention Prediction
The module of human intention prediction determines current human actions and aggregates them into a history record. Here, we implement an algorithm that leverages Bayesian networks to fuse three modalities of human information: head orientation, hand orientation, and hand movement [24]. By fusing three modalities for human intention recognition, the accuracy and precision of the prediction are dramatically increased because different modality provides valuable information in different stages of the actions.
@@ III-E Relevance Quantification with LLMs
The first step in quantifying relevance with LLM is contextualizing the scene. Currently, this is achieved with a fixed prompt $\mathcal{P}$ designed and optimized specifically for the setup and the problem of HRC. However, in the future, automatic scene contextualization and prompt construction can be considered and developed. In the prompt $\mathcal{P}$, the environment env, the objects in the scene object in env, and the action history of the human action history are automatically generated from the scene understanding and human intention prediction modules. An example of the prompt is: In a kitchen with [object in env], a person has already grabbed [action history]. What objective is the human trying to finish? What objects will be relevant next among [object in env]? In the prompt, [object in env] and [action history] can be derived and populated automatically. The output format regulation of the prompt is omitted due to the page limit.
The outputs of the LLM module are the prediction of the human objective and the set of relevant elements $\mathcal{E}_{r}$. With this method, world knowledge of an LLM can be extracted for a general environment and setup without any training and transfer learning.
@@ III-F Decision Making - Task Allocation
In the problem definition, we define the subscript $j$ representing the index of elements such that $e_{j}\in\mathcal{E}_{r}$. We define the position vector of element $e_{j}$ in the world coordinate as $\mathbf{p}_{e_{j}}$, the position of the common destination of relevant elements as $\mathbf{p}_{d}$, the initial location of the robot as $\mathbf{p}_{r_{0}}$, the velocity of the robot as $v_{r}$, the velocity of human as $v_{h}$, the start-up delay for the human to finish the current task as $d_{h}$. We consider $d_{h}$ because there is a temporal delay between the time the robot and the human start to fulfill the allocated tasks. We further define $T_{r}$ and $T_{h}$ as the time durations for the robot or the human to finish the assigned tasks, which can be computed as:
@@ III-G Decision Making - Motion Planning
where $A$ is the magnitude of the attractive force, $d_{g}$ is the distance between the robot and its goal, $\alpha_{a}$ is the constant that controls how the attractive force increases with distance, and $\mathbf{\hat{u}_{a}}$ is the unit vector from the robot to the goal.
@@
To consider the future motion of the human hand, we developed a novel methodology to construct an ellipsoid virtual obstacle, as shown in Fig. 2. The two endpoints of the major axis of the ellipsoid are the current position of the human hand $\mathbf{p}_{h}$ and the position of the predicted hand destination $\mathbf{p}_{t}$. The semi-axes $a$, $b$, and $c$ of the ellipsoid are given by:
@@
where $\mathbf{\hat{u}_{h\rightarrow t}}$ represents the unit vector from $\mathbf{p}_{h}$ to $\mathbf{p}_{t}$, and $t_{s}$ is a time factor we proposed in the unit of $s$ reflecting the available safety buffer in terms of time. $k_{r}$ decreases as $\mathbf{p}_{c}$ becomes farther away from $\mathbf{p}_{h}$. Without $k_{r}$, $\mathbf{F}_{r_{v}}$ will consider the virtual obstacle too early and conservatively, resulting in unnecessary detours and a longer robot trajectory.
@@
This total force determines the direction and magnitude of the robot’s motion, allowing it to avoid both physical and virtual obstacles while moving toward its goal.
@@ IV-A Dataset
The performance of human objective prediction and relevance prediction is assessed via the Breakfast Actions Dataset [26], which comprises a variety of typical human activities performed during breakfast time (e.g., preparing coffee, cooking pancakes, making hot chocolate, etc.). For each test, ground truth data provides the ground truth objective (GTO), and the ground truth plan (GTP), which describes the sequence of actions performed by an individual in the execution of the objective. In our evaluation, a step ratio of segments from the GTP is contextualized and fed into LLM to predict the human objective and the set of relevant elements. Three distinct ratio step values, 0.25, 0.5, and 0.75, are employed. The objective prediction is evaluated through a manual assessment process. The predicted relevant objects are evaluated automatically with word matching. The metrics for the evaluation are percentages of tests with correct prediction of the objective and the relevant elements.
@@ IV-B Simulation Development and Setup
To verify the effectiveness of relevance for safer HRC, we developed a simulator as shown in Fig. 3. A large number of simulations can be carried out systematically and identically for different test cases, facilitating the evaluation of the decision making module. In the simulator, we constructed the environment with a table (the cyan cuboid) and objects on top of the table. The height of the table is set to be 73 cm. The table size is 180 cm in the $x$ direction, 6 cm in the $z$ direction, and 76 cm in the $y$ direction. The objects on the table consist of two parts: the necessary objects for the objective and randomly added objects from a list of kitchen objects. The locations for each object on the tabletop are randomly selected from a collection of locations uniformly distributed on a half circle with a radius of 60 cm. This object arrangement enables a wide range of relative velocities and relative angles between moving objects during testing, recovering realistic scenarios. The diameter of each object is 8 cm.
@@ V-A Accurate Objective and Relevance Prediction
TABLE I: Performance of objective prediction and relevance prediction for the 10 objectives in the Breakfast Actions Dataset with various step ratios of action as inputs. These results show that our methodology can accurately and robustly predict human objectives and relevance.
@@
We first analyze the performance of objective prediction using our methodology, upon which relevance determination depends. Upon further examination of individual objectives, certain objectives, such as the preparation of cereal, coffee, hot chocolate, juice, salad, and tea, exhibit consistently high predictive accuracy across all evaluated ratio steps. The prediction accuracy for those objectives is high even at a low step ratio of 0.25, demonstrating our methodology can accurately identify the human’s objective at an early stage of the human action and motion. Consistent with expectations, an increase in the “ratio steps” parameter correlates positively with enhanced prediction accuracy. The contextualization of the scene with a higher step ratio will contain more indicative cues about the objectives and thus improve the objective prediction accuracy. The inaccuracy for fried eggs is attributed to the inherent similarity between the objectives involving eggs, such as omelettes and scrambled eggs. However, the relevance for those objectives is very similar.
Next, we analyze the performance of relevance prediction using our methodology based on the predicted objective. The relevance prediction performs extraordinarily for specific objectives, including hot chocolate, juice, and salad, which achieves a 100% accuracy at the step ratio of 0.25. The prediction accuracy of all other objectives achieves at least 0.85 at the step ratio of 0.75. For the relevance prediction, the accuracy also increases with the step ratio. It’s worth noting that the relevance prediction for the objectives predicted with low accuracy can still be accurate. This is attributed to the fact that similar objectives, such as omelettes and scrambled eggs, share the same relevance, which bolsters the relevance prediction despite inaccuracies in objective identification.